# Large Language Models as Urban Residents: An LLM Agent Framework for Personal Mobility Generation

### v1

@@ 4.2 Motivation-Driven Activity Generation
In Phase 2 of LLMob, we focus on the retrieval of motivation and the integration of motivation and activity patterns for individual activity trajectory generation. Retrieval augmentation has been identified as a crucial factor in boosting the performance of LLM Xu et al. (2023). This enhancement provides additional information that aids LLM in more effectively responding to queries. While previous studies on activity generation mainly overlook the critical factors of macro temporal information (i.e., date) or specific scenarios (i.e., hash weather) Yuan et al. (2022), we propose a more sophisticated activity generation which accounts for various conditions by taking advantage of the human-like intelligence of LLM. For instance, the activity trajectory at date $d$ can be inferred given the motivation of this date and the normal activity pattern as:
@@
This activity generation scheme instructs the LLM to simulate a designated individual according to a given activity pattern, and then meticulously generate an activity trajectory in accordance with the daily motivation. To obtain insightful and reliable motivations for the generation, two retrieval schemes are proposed.
@@ 4.2.1 Evolving-based Motivation Retrieval
This scheme is related to the intuitive principle that an individual’s motivation on any given day is influenced by her interests and priorities in preceding days Park et al. (2023). Guided by this understanding that new motivations often evolve from existing ones, our approach harnesses the intelligence of LLM to understand the behavior of daily activities and the underlying motivations. As illustrated in Fig. 3, for a specific day $d$ for which we aim to generate the activity trajectory, as well as considering the activities of the past few days, we prompt the LLM to act as an individual based on the pattern identified in Section 4.1 and summarize the motivation behind these activities. Leveraging these identified past motivations, the LLM is further prompted to infer potential motivations for the subsequent day.
Figure 3: Evolving-based motivation retrieval.
@@ 4.2.2 Learning-based Motivation Retrieval
In this scheme, we hypothesize that individuals tend to establish routines in their daily activities, guided by consistent motivations even if the specific locations may vary. For example, if someone frequently visits a burger shop on weekday mornings, this behavior might suggest a motivation for a quick breakfast. Based on this, it is plausible to predict that the same individual might choose a different fast food restaurant in the future, motivated by a similar desire for convenience and speed during their morning meal. We introduce a learning-based scheme to retrieve motivation from historical data. For each new date on which to plan activities, the only information available is the date itself. To use this clue for planning, we first formulate a relative temporal feature $\boldsymbol{z}_{{d_{c},d_{p}}}$ between a past date $d_{p}$ and the current date $d_{c}$. This feature captures various aspects, such as the gap between these two dates and whether they belong to the same month. Utilizing this setting, we train a score approximator $f_{\theta}(\boldsymbol{z}_{{d_{c},d_{p}}})$ to evaluate the similarity between any two dates. Notably, due to the lack of supervised signals, we employ unsupervised learning to train $f_{\theta}(\cdot)$. Particularly, a learning scheme based on contrastive learning Chen et al. (2020) is established. For each trajectory of a user, we can scan her other trajectories and identify similar (positive) and dissimilar (negative) dates according to a predefined similarity score. This similarity score is calculated between two activity trajectories $\mathcal{T}_{d_{a}}$ and $\mathcal{T}_{d_{b}}$ as:
$sim_{d_{a},d_{b}}=\sum_{t=1}^{N_{d}}\mathbf{1}_{(\mathcal{T}_{d_{a}}(t)=\mathcal{T}_{d_{b}}(t))}\text{ if }|\mathcal{T}_{d_{a}}|>t\text{ and }|\mathcal{T}_{d_{b}}|>t,$ (3)
where $N_{d}$ is the total number of time intervals (e.g., 10 min) in one day. $\mathcal{T}_{d_{a}}(t)$ indicates the $t$th visiting location recorded in trajectory $\mathcal{T}_{d_{a}}$. Intuitively, there should be more shared locations in the similar trajectory pair. Thereafter, positive pair is characterized by the highest similarity score, indicative of a greater degree of resemblance between the trajectories. Conversely, negative pairs are marked by low similarity scores, reflecting a lesser degree of commonality. After obtaining the training dataset from these positive and negative pairs, we can train a model to approximate the similarity score between any two dates through contrastive learning. Specifically, the procedure involves the following steps:
@@
    Adopt InfoNCE Oord et al. (2018) as the contrastive loss function:
    $\mathcal{L}(\theta)=\sum_{n=1}^{N}-\log\left(\frac{e^{\text{logits}_{i}}}{\sum^{k+1}_{j=1}e^{\text{logits}_{j}}}\right)_{n},$ (5)
    where $N$ is the batch size of the samples, $i$ indicates the index of the positive pair.
Upon training a similarity score approximation model, it can be applied to access the similarity between any given query date and historical dates. This enables us to retrieve the most similar historical data, which allows the LLM to generate a summary of the motivations prevalent at that time. By doing so, we can extrapolate a motivation relevant to the query date, providing a basis for the LLM to generate a new activity trajectory.
5 Experiment
@@ 5.1 Experimental Setup
Dataset. We investigate and validate LLMob over a personal activity trajectory dataset from Tokyo. This data set is obtained through the Twitter API and the Foursquare API, which covers data from January 2019 to December 2022.
Table 1: Sample of the personal activity data.
UserID Latitude Longitude Location Name Category Time 12199425 35.508 139.615 Convenience Store Shop & Service 8:00 12199425 35.509 139.618 Seafood Restaurant Food & Service 8:30
@@
The time frame of this data set is insightful as it captures typical daily life prior to the COVID-19 pandemic (i.e., normal period) and subsequent alterations during the pandemic (i.e., abnormal period). To facilitate an efficient and detailed analysis for different periods, we randomly choose 100 users to model their individual activity trajectory at a 10-minute interval according to the number of available trajectories (sample given in Table 1). We utilize the category classification in Foursquare to determine the activity category for each location. We use 10 candidate personas as a prior for subsequent pattern generation, as shown in Table 2, which adequately capture a diverse range of activity patterns within the data of this study. For the application to other datasets, this style of candidate patterns can be easily initialized using an LLM.
Table 2: Suggested personas and corresponding detailed description.

### v2

@@ 3.2 Motivation-Driven Activity Generation
In Phase 2 of LLMob, we focus on the retrieval of motivation and the integration of motivation and activity patterns for individual activity trajectory generation. Retrieval-augmented generation has been identified as a crucial factor in boosting the performance of LLM [35]. This enhancement provides additional information that aids LLM in more effectively responding to queries. While previous studies on activity generation mainly overlook the critical factors of macro temporal information (e.g., date) or specific scenarios (e.g., harsh weather) [37], we propose a more sophisticated activity generation which accounts for various conditions by taking advantage of the human-like intelligence of LLM. For instance, the activity trajectory at date $d$ can be inferred given the motivation of this date and the habitual activity pattern as:
@@
This generation scheme instructs the LLM agent to simulate a designated individual according to a given activity pattern, and then meticulously generate an activity trajectory in accordance with the daily motivation. To obtain insightful and reliable motivations, two retrieval schemes are proposed.
Figure 3: Evolving-based motivation retrieval.
@@ 3.2.1 Evolving-based Motivation Retrieval
This scheme is related to the intuitive principle that an individual’s motivation on any given day is influenced by her interests and priorities in preceding days [27]. Guided by this understanding, our approach harnesses the intelligence of the LLM agent to understand the behavior of daily activities and the underlying motivations. As illustrated in Figure 3, for a specific date $d$ for which we aim to generate the activity trajectory, we consider the activities of the past $k$ days ($k=\min(7,l)$, where $l$ is the maximum value such that the trajectory for date $d-l$ can be found in the database), and prompt the LLM agent to act as an urban resident based on the pattern identified in Section 3.1 and summarize $k$ motivations behind these activities. Leveraging these summarized motivations, the LLM agent is further prompted to infer potential motivation for date $d$.
@@ 3.2.2 Learning-based Motivation Retrieval
In this scheme, we hypothesize that individuals tend to establish routines in their daily activities, guided by consistent motivations even if the specific locations may vary. For example, if someone frequently visits a burger shop on weekday mornings, this behavior might suggest a motivation for a quick breakfast. Based on this, it is plausible to predict that the same individual might choose a different fast food restaurant in the future, motivated by a similar desire for convenience and speed during their morning meal. We introduce a learning-based scheme to retrieve motivation from historical data. For each new date on which to plan activities, the only information available is the date itself. To use this clue for planning, we first formulate a relative temporal feature $\boldsymbol{z}_{{d_{c},d_{p}}}$ between a past date $d_{p}$ and the current date $d_{c}$. This feature captures various aspects, such as the gap between these two dates and whether they belong to the same month. Utilizing this setting, we train a score approximator $f_{\theta}(\boldsymbol{z}_{{d_{c},d_{p}}})$ to evaluate the similarity between any two dates. Notably, due to the lack of supervised signals, we employ unsupervised learning to train $f_{\theta}(\cdot)$. Particularly, a learning scheme based on contrastive learning [5] is established. For each trajectory of a resident, we can scan her other trajectories and identify similar (positive) and dissimilar (negative) dates according to a predefined similarity score. This similarity score is calculated between two activity trajectories $\mathcal{T}_{d_{a}}$ and $\mathcal{T}_{d_{b}}$ as:
$sim_{d_{a},d_{b}}=\sum_{t=1}^{N_{d}}\mathbf{1}_{(\mathcal{T}_{d_{a}}(t)=\mathcal{T}_{d_{b}}(t))}\text{ if }|\mathcal{T}_{d_{a}}|>t\text{ and }|\mathcal{T}_{d_{b}}|>t$ (3)
where $N_{d}$ is the total number of time intervals (e.g., 10 min) in one day. $\mathcal{T}_{d_{a}}(t)$ indicates the $t$th visiting location recorded in trajectory $\mathcal{T}_{d_{a}}$. Intuitively, there should be more shared locations in the similar trajectory pair. Thereafter, the positive pair is characterized by the highest similarity score, indicative of a greater degree of resemblance between the trajectories. Conversely, the negative pairs are marked by low similarity scores, reflecting a lesser degree of commonality. After obtaining the training dataset from these positive and negative pairs, we train a model to approximate the similarity score between any two dates by contrastive learning. This procedure involves the following steps:
@@
    Adopt InfoNCE [24] as the contrastive loss function:
    $\mathcal{L}(\theta)=\sum_{n=1}^{N}-\log\left(\frac{e^{\text{logits}_{i}}}{\sum^{k+1}_{j=1}e^{\text{logits}_{j}}}\right)_{n}$ (5)
    where $N$ is the batch size of the samples and $i$ indicates the index of the positive pair.
Upon training a similarity score approximation model, it can be applied to access the similarity between any given query date and historical dates. This enables us to retrieve the most similar historical data, which is prompted to the LLM agent to generate a summary of the motivations prevalent at that time. By doing so, we can extrapolate a motivation relevant to the query date, providing a basis for the LLM agent to generate a new activity trajectory.
4 Experiments
@@ 4.1 Experimental Setup
Dataset. We investigate and validate LLMob over a personal activity trajectory dataset from Tokyo. This dataset was obtained through Twitter and Foursquare APIs and covers the data from January 2019 to December 2022. The time frame of this dataset is insightful as it captures typical daily life prior to the COVID-19 pandemic (i.e., normal period) and subsequent alterations during the pandemic (i.e., abnormal period). To facilitate a cost-efficient and detailed analysis for different periods, we randomly choose 100 users to model their individual activity trajectory at a 10-minute interval according to the number of available trajectories (samples given in Appendix C.2). We utilize the category classification in Foursquare to determine the activity category for each location. We use 10 candidate personas (Appendix C.4) as a prior for subsequent pattern generation, which captures a diverse range of activity patterns within the data of this study. For the application to other datasets, this style of candidate patterns can be easily initialized using an LLM.
Metrics. The following features related to personal activity are used as metrics: (1) Step distance (SD)[37]: This metric evaluates the spatial pattern of an individual’s activities by measuring the distance between two consecutive locations in a trajectory. (2) Step interval (SI)[37]: This metric evaluates the temporal pattern of an individual’s activities by measuring the time interval between two successive locations on an individual’s trajectory. (3) Daily activity routine distribution (DARD): This metric captures and examines the distribution of activities for each trajectory, focusing on the patterns of individual activities characterized by activity type and timing (e.g., activity type, time). It provides insight into how activities are distributed over space and time and reflects semantic information such as habitual behavior. (4) Spatial-temporal visits distribution (STVD): This metric provides a granular perspective on the generated activities by assessing the spatial-temporal distribution of visited locations within each trajectory, including geographical coordinates and timestamps. It enables a detailed analysis of where and when activities occur.
After extracting the above features from both the generated and real trajectory data, Jensen-Shannon divergence (JSD) is employed to quantify the discrepancy between them. Lower JSD is preferred.
Methods. LLMob is evaluated against: Markov-based mechanic model (MM) [26], an LSTM-based prediction model (LSTM) [12], two attention-based prediction models, including DeepMove [8] and STAN [21]. Within the domain of deep generative models, we select two adversarial learning frameworks, including TrajGAIL [6] and ActSTD [37], as well as a diffusion model, DiffTraj [41]. We use the source codes of the baselines provided by their respective authors and adapt them to our setting.
[…]
4.2 Main Results and Analysis
Generative Performance Validation (RQ 1, RQ 2). The performance evaluation involves analyzing generation results in three distinct settings: (1) Generating normal trajectories based on normal historical trajectories in 2019, a period unaffected by the pandemic. (2) Generating abnormal trajectories based on abnormal historical trajectories in 2020, a year marked by the pandemic. (3) Generating abnormal trajectories in 2021 (pandemic) based on normal historical trajectories in 2019.
The results of these evaluations are detailed in the metrics reported in Table 1. Through the comparison, it can be observed that although LLMob may not excel in replicating spatial features (SD) precisely, it demonstrates superior performance in handling temporal aspects (DI). When considering spatial-temporal features (DARD and STVD), LLMob’s performance is also competitive. In particular, LLMob achieves the best performance on DI and DARD for all the three settings, and is the runner-up on STVD. Baselines like DeepMove and TrajGAIL perform the best on SD and STVD, respectively, but become much less competitive when evaluated in other aspects. We suggest that the pronounced advantage of LLMob in terms of DARD (roughly 1/2 to 1/3 JSD compared to the best of baselines) can be attributed to the LLM agent’s tendency to accurately replicate the motivation behind individual activity behaviors. For instance, an agent may recognize patterns like a person’s habits to have breakfast in the morning, without being restricted to a specific restaurant. This phenomenon highlights the enhanced semantic understanding capabilities of the LLM agent.
Table 1: Performance (JSD) of trajectory generation based on historical data. Lower is better. Winners and runners-up are marked in boldface and underline, respectively.
Models Normal Trajectory, Normal Data Abnormal Trajectory, Abnormal Data Abnormal Trajectory, Normal Data
(# Generated Trajectories: 1497) (# Generated Trajectories: 904) (# Generated Trajectories: 3555)
SD SI DARD STVD SD SI DARD STVD SD SI DARD STVD
MM [26] 0.018 0.276 0.644 0.681 0.041 0.300 0.629 0.682 0.039 0.307 0.644 0.681
LSTM [12] 0.017 0.271 0.585 0.652 0.016 0.286 0.563 0.655 0.035 0.282 0.585 0.653
DeepMove [8] 0.008 0.153 0.534 0.623 0.011 0.173 0.548 0.668 0.013 0.173 0.534 0.623
STAN [21] 0.152 0.400 0.692 0.692 0.115 0.092 0.693 0.691 0.142 0.094 0.692 0.690
TrajGAIL [6] 0.128 0.058 0.598 0.489 0.133 0.060 0.604 0.523 0.332 0.058 0.434 0.428
ActSTD [37] 0.034 0.436 0.693 0.692 0.071 0.469 0.692 0.692 0.022 0.093 0.468 0.692
DiffTraj [41] 0.052 0.251 0.318 0.692 0.008 0.240 0.339 0.692 0.101 0.142 0.218 0.693
LLMob-E 0.053 0.046 0.125 0.559 0.056 0.043 0.127 0.615 0.062 0.056 0.117 0.536
LLMob-E w/o $\mathcal{P}$ 0.055 0.069 0.223 0.530 0.059 0.081 0.252 0.673 0.065 0.079 0.209 0.561
LLMob-E w/o $\mathcal{SC}$ 0.058 0.076 0.295 0.589 0.068 0.086 0.225 0.649 0.072 0.096 0.301 0.589
LLMob-L 0.049 0.054 0.136 0.570 0.057 0.051 0.124 0.609 0.064 0.051 0.124 0.531
LLMob-L w/o $\mathcal{P}$ 0.061 0.080 0.270 0.600 0.072 0.081 0.286 0.641 0.073 0.091 0.248 0.580
LLMob-L w/o $\mathcal{SC}$ 0.057 0.074 0.236 0.602 0.071 0.084 0.236 0.642 0.073 0.094 0.286 0.622
LLMob w/o $\mathcal{M}$ 0.059 0.078 0.264 0.590 0.066 0.080 0.274 0.633 0.074 0.090 0.255 0.563
[…]
Exploring Utility in Real-World Applications (RQ 3). We are interested in how LLMob can elevate the social benefits, particularly in the context of urban mobility. To this end, we propose an example of leveraging the flexibility and intelligence of LLM agents in understanding semantic information and simulating an unseen scenario. In particular, we enhance the original setup by incorporating an additional prompt to provide a context for the LLM agent, enabling it to plan activities during specific circumstances. For example, a “pandemic” prompt is as follows: Now it is the pandemic period. The government has asked residents to postpone travel and events and to telecommute as much as possible.
Figure 4: Daily activity frequency.
(a) Arts & entertainment.
(b) Professional & other places.
Figure 5: Activity heatmaps for the pandemic scenario.
By integrating the above prompt, we can observe the impact of external elements, such as the pandemic and the government’s measures, on urban mobility and related social dynamics. We use the activity trajectory data during the pandemic (2021) as ground truth and plot the daily activity frequency in 7 categories in Figure 4. TrajGAIL, despite delivering the best STVD in Table 1, displays very low frequencies for all the categories, and fails to reflect the tendency of each category. In contrast, a comparison between LLMob-L and the one augmented with the pandemic prompt demonstrates the impact of external factors: there is a significant decrease in activity frequency with the pandemic prompt, which semantically discourages activities likely to spread the disease (e.g., food).
Additionally, from a spatial-temporal perspective, two major activities (e.g., Arts & entertainment and Professional & other places) are selected to observe the behavior, as shown in Figures 5(a) and  5(b). These activities are particularly insightful as they encapsulate the impact of the pandemic on the work-life balance and daily routines of residents. Specifically, with the pandemic prompt, LLMob reproduces a more realistic spatial-temporal activity pattern. This enhanced realism in the generation is attributed to the integration of prior knowledge about the pandemic’s effects and governmental responses, allowing the LLM agent to behave in a manner that aligns with actual behavioral adaptations. For instance, the reduction in Arts & entertainment activities reflects the closure of venues and social distancing guidelines, while changes in Professional & other places activities indicate shifts toward remote work and the transformation of professional environments. Intuitively, prompting the LLM agent to generate activities based on various priors shows great potential in real-world applications. The utility of such a conditioned generative approach, coupled with the reliable generated results, can significantly alleviate the workload of urban managers. We suggest that this kind of workflow can simplify the analysis of urban dynamics and aid in assessing the potential impact of urban policies.
4.3 Ablation Studies
Impact of Patterns. In Table 1, by comparing LLMob with and without using patterns (“w/o $\mathcal{P}$”), we observe that the identified patterns consistently enhance the trajectory generation performance. The improvement on DARD is the most significant (reducing JSD by around 50%), showcasing the use of patterns is a key factor in capturing the semantics of daily activity. We provide example patterns in Appendix D.1 to show how the habitual behaviors of individuals are recognized by patterns.
Impact of Self-Consistency Evaluation. By comparing LLMob with and without self-consistency evaluation (“w/o $\mathcal{SC}$”) in Table 1, we find that self-consistency is useful in all aspects, and its impact is the most significant on DARD, especially when generating abnormal trajectories from normal data, showcasing its effectiveness in processing semantics. We also observe that “w/o $\mathcal{SC}$” performs even worse than “w/o $\mathcal{P}$” in many cases, because in “w/o $\mathcal{SC}$”, a candidate pattern is randomly picked for summarizing motivations, potentially introducing inconsistency to an individual’s daily activity.
[…]
Impact of Motivation Retrieval Strategy. We compare LLMob equipped with the two motivation retrieval strategies (“-E” and “-L”). Table 1 shows that no retrieval strategy always outperforms the other, though evolving-based retrieval wins in more cases (7 vs 5). Moreover, evolving-based retrieval is generally less sensitive to the removal of patterns or self-consistency evaluation, suggesting that resorting to the LLM to process historical trajectories is more robust than using contrastive learning.
5 Conclusion
Contributions. This study is believed to be the first personal mobility simulation empowered by LLM agents on real-world data. Our innovative framework leverages activity patterns and motivations to direct LLM agents in emulating urban residents, facilitating the generation of interpretable and effective individual activity trajectory. Extensive experimental studies based on real-world data are conducted to validate the proposed framework and demonstrate the promising capabilities of LLM agents to improve urban mobility analysis.
Social Impacts. Leveraging artificial intelligence to enhance societal benefits is increasingly promising, especially with the advent of high-capacity models such as LLMs. This study explores one of the potential avenues for applications using LLMs as reliable agents to simulate specific scenarios to assess the effects of external factors, such as pandemic and government policies. The introduced framework offers a flexible approach to enhance the reliability of LLMs in simulating urban mobility.
Limitations. In this study, we considered individual agents’ activities without interactions between them. We will target the multi-agent case to model interactions (e.g., people may follow others in a friends’ or family’s activity) as future work. Due to the difficulty in collecting high-quality personal mobility data (most datasets are incomplete in showing one’s daily activity), only one dataset is used in the experimental evaluation. Moreover, only GPT-3.5 was evaluated for the sake of cost-efficiency, whereas more advanced LLMs may yield better generation performance.
Appendix
Appendix A Algorithm Pseudo-Codes
The pseudo-code of the algorithm for activity pattern identification is given in Algorithm 1.
Input : Number of personas $C$, activity trajectories $\mathcal{T}$.
Output : Best activity pattern $best\_pattern$.
1
2Initialize: Empty candidate pattern set $\mathcal{CP}$;
4 for $i=1$ to $C$ do
5       Formulate prompt $p_{1}$ with prior information and persona $i$;
6       $cp\leftarrow\text{LLM}(p_{1})$;
7       $\mathcal{CP}\leftarrow\mathcal{CP}\cup\{\,cp\,\}$;
8
9
10Initialize a score dictionary $\mathcal{S}$ to store pattern scores;
11
12foreach candidate pattern $cp$ in $\mathcal{CP}$ do
13       Initialize $score_{cp}=0$;
14       foreach activity trajectory $t$ in $\mathcal{T}$ do
15             Formulate evaluation prompt $p_{2}$ for $cp$ and $t$;
16             $score_{cp}\leftarrow score_{cp}+\text{LLM}(p_{2})$;
17      $\mathcal{S}[cp]\leftarrow score_{cp}$;
18
19
[…]
return $best\_pattern$
Algorithm 1 Activity Pattern Identification
Appendix B Prompts
where $<$INPUT 0$>$ and $<$INPUT 1$>$ are replaced by the candidate personas, and $<$INPUT 2$>$ and $<$INPUT 3$>$ will be replaced by the activity habits extracted from historical data. Specifically, we formulate the $<$INPUT 2$>$ in the following format:
$<$INPUT 2$>$ format During $<$weekend or weekday$>$, you usually travel over $<$distance$>$ kilometers a day, you usually begin your daily trip at $<$time of first daily activity$>$ and end your daily trip at $<$time of last daily activity$>$. you usually visit $<$location of first daily activity$>$ at the beginning of the day and go to $<$location of last daily activity$>$ before returning home.
where $<$INPUT 0$>$ is the candidate pattern, and $<$INPUT 1$>$ is the daily activities plan for evaluation.
where $<$INPUT 0$>$ is replaced by the selected pattern, $<$INPUT 1$>$ is replaced by the number of days from the date to plan activities, and $<$INPUT 2$>$ is the historical activities corresponding to the chosen date.
where $<$INPUT 0$>$ is replaced by the selected pattern, and $<$INPUT 1$>$ Is replaced by the retrieved historical activities.
[…]
#125 at 9:10, Town Hall
#489 at 12:50, Rest Area
#585 at 13:40, Seafood Restaurant
#105 at 14:20] “reason”: “My plan today is to finish my teaching duty in the morning and find something delicious to taste.”}}\}}
where $<$INPUT 0$>$ is replaced by the selected pattern, $<$INPUT 1$>$ is replaced by the retrieved motivation, and $<$INPUT 2$>$ is replaced by the most frequently visited locations.
Pandemic Now it is the pandemic period. The government has asked residents to postpone travel and events and to telecommute as much as possible.
Appendix C Experimental Setup
C.1 Environment
We leverage the GPT API to conduct our generation studies. Specifically, we use the gpt-3.5-turbo-0613 version of the API, which is a snapshot of GPT-3.5-turbo from June 13th, 2023. The experiments were carried out on a server with the following specifications:
  • CPU: AMD EPYC 7702P 64-Core Processor
    • Architecture: x86_64
    • Cores/Threads: 64 cores, 128 threads
    • Base Frequency: 2000.098 MHz
  • Memory: 503 GB
  • GPUs: 4 x NVIDIA RTX A6000
    • Memory: 48GB each
C.2 Personal Activity Dataset
We use a personal activity trajectory dataset from Tokyo which was obtained through Twitter and Foursquare APIs, covering the data from January 2019 to December 2022. We randomly choose 100 users to model their individual activity trajectory at a 10-minute interval according to the number of available trajectories. Samples are shown in Table 2.
Table 2: Samples of personal activity data.
UserID Latitude Longitude Location Name Category Time
44673 35.008 139.015 Convenience Store Shop & Service 2019-12-17 8:00
44673 35.009 139.018 Ramen Restaurant Food & Service 2019-12-17 8:30
44673 35.004 139.060 Italian Restaurant Food & Service 2019-12-17 11:20
44673 35.009 139.085 Farmers Market Shop & Service 2019-12-17 14:20
44673 35.005 139.086 Soba Restaurant Food & Service 2019-12-17 18:00
C.3 Learning-Based Motivation Retrieval
Table 3: Architecture of the score approximator.
Layer Input Size Output Size Notes
Input Layer 3 64 Linear
Activation - - ReLU
Output Layer 64 1 Linear
We include the day of the year for the query date, whether it shares the same weekday as the reference date, and whether both the query and reference dates fall within the same month as input features. Settings for the learning process are as follows: Adam [16] is used as the optimizer, batch size is 64, learning rate is 0.002, and the number of negative samples is 2.
C.4 Personas
We use 10 candidate personas as a prior infromation for subsequent pattern generation, as shown in Table 4.
Table 4: Suggested personas and corresponding descriptions.